Science and medicine. A plane flies 720 mi against a steady 30 mi/h headwind and then returns to the same point with the wind. If the entire takes 10 h, what is the plane's speed in still air?
Answer by ptaylor(1894)


Distance(d) equals rate(r) times time(t) or d=rt; t=d/r and r=d/t
Let r=plane's speed in still air
Now we are told that time traveled against the wind plus time traveled with the wind equals 10h
Time traveled against the wind=720/(r-30)
Time traveled with the wind=720/(r+30)
So:
720/(r-30)+720/(r+30)=10 Multiply each term by (r+30)(r-30) to get rid of fractions
720(r+30)+720(r-30)=10(r+30)(r-30) simplify
720r+30(720)+720r-30(720)=10(r^2-900) or
1440r=10r^2-9000 subtract 1440r from both sides and divide both sides by 10:
r^2-144r-900=0 quadratic in standard form and this can be factored:
(r-150)(r+6)=0
r=150 mph-------------------------------------answer
and
r=-6 mph---------------discount negative speeds
CK
720/(150-30)+720/(150+30)=10
720/120+720/180=10
6+4=10
10=10

Hope this helps------ptaylor